### v1

@@ 5.1 Dataset
To fine-tune and evaluate PerfRL, we use the dataset from PIE Madaan et al. (2023). This dataset consists of approximately 40k Python files, 88k C++ files, and 3.6k Java files. PIE captures the progressive changes made by a programmer over time to improve their code. The dataset also contains (slow, fast) pairs of code written by the same programmer. We run our experiments on the subset of the dataset that concerns the Python files. The training, test and validation sets consist of approximately 36k, 1k and 2k samples respectively. Each of the samples has at least one associated unit test file that requires a specific input and has an expected output result. We test the accuracy of the dataset by executing all the input source code with a 5-second execution limit. We found that the $72.4\%$ of the training data, $76.4\%$ of the testing data, and $70.8\%$ of the validation data are executable. During our training, we skip all the input code that is not originally executable in the first place. As for testing and validation, we use all the data in order to compare our approach with the baseline model.
5.2 Setup
@@
We run the fine-tuning and reinforcement learning steps for an instance of the CodeT5 model with $60$ million parameters and a learning rate of $2\times 10^{-5}$. To reduce the training time we run the whole process for $8$ RL steps for $3$ epochs. For the training of the model, we use an NVIDIA A100 GPU graphics card with 40GB of RAM on an Ubuntu 20.04 server with 50 cores. Our model is trained within approximately 30 hours. Each epoch computes (i) the score of the generated sequence (see Equation 9) and (ii) the loss for all the samples of the dataset from the same input data. During the training phase, we set the temperature to 1 and top_k to 50 for random sampling.
@@
As already mentioned, prior to running the reinforcement learning step, we fine-tune the CodeT5 model with a one-shot learning setting. That is to say, we feed each sample of the dataset into the model once and compute the loss as described in Equation 3. We set the learning rate to $5\times 10^{-5}$ and batch size to $32$. The results are available in Table 1.
@@
Table 1: Evaluation results of PerfRL and baseline models. The bold font indicates the evaluation results of PerfRL. The asterisk (*) indicates the baseline model. The first block shows the results as reported in the paper of PIE Madaan et al. (2023)
@@
METHOD Model Size Sample strategy %OPT SP RTR
CODEGEN-16B 16B greedy and 1-shot 2.2 1.55 25.05
CODEGEN-2B 2B greedy 8.2 2.32 48.23
CODEGEN-16B 16B greedy 14.6 1.69 51.25
CODEX greedy 14.3 2.7 53.49
CodeT5 (Before RL) 60M greedy and 0-shot 0 0 0
CodeT5 (24 Fine-tuning epoch)* 60M greedy and 0-shot 0.5 2.27 53.42
PerfRL (8 RL steps) 60M greedy and 0-shot 2.8 4.93 36.93
@@
For our experiments, we set the reward values for the reward function $r$ to $R1=0$, $R2=1$, $R3=1.3$ and $R4=2$. As already mentioned in Section 4.3.1, the actual reward values do not affect the performance of the RL step, as long as we assign a higher value to a much preferable sample.
5.3 Baselines
We use a fine-tuned version of the CodeT5 model with $24$ epochs as our main baseline. Furthermore, we compare our results with those of the original PIE paper, where models from the family of CodeGen and Codex are used for their experiments. We benchmark our approach against the main baseline model using the evaluation metrics described in Section 5.4.
5.4 Evaluation Metrics
We use the following evaluation metrics as also defined in the paper of the PIE Madaan et al. (2023) dataset:
  • Percent Optimized (%OPT): The ratio of samples on the test set that are improved by a given method.
@@ 5.4 Evaluation Metrics
To measure the execution time of each generated program, we calculate the cumulative execution time using all the unit tests. We run each experiment three times and measure the mean time on each set of unit tests in order to ensure that the reduction of the execution time is not affected by random factors.
5.5 Evaluation Results
@@
Table 1 presents the mean %OPT, SP, and RTR scores of the proposed method as well as that of the baseline model. Although PerfRL relies on a model with fewer parameters, it manages to perform equally or even better with respect to the greedy and 1-shot versions of the CodeGen-16B and CodeGen-2B models. Furthermore, as illustrated in Figure 3, both the pass and optimization rate increase proportional to the RL steps. Thus, we argue that the utilization of an RL step enables smaller models to learn easily the task of source code optimization in a less resource and time-consuming manner. More specifically, we train our model for 30 hours as opposed to the baseline CodeGen which is trained for approximately 3 days and requires a stronger machine (i.e. 2 $\times$ NVIDIA A6000 for CodeGen-2B and 4 $\times$ NVIDIA A6000 for CodeGen-16B).
@@
In order to determine whether our RL step is able to generate good candidates by random or greedy sampling during the training, we measure the compilation, pass, and optimization rate as shown in Figure 4. Considering our sampling strategy that concatenates the target program with the generated samples, we empirically observe that the threshold for these three values is approximately $20\%$ (red dashed line). Since these three rates are always above the threshold, they denote that our model generates meaningful candidates by itself other than purely based on the target program.
Figure 3: Pass rate and optimization rate on validation data over RL steps on the fine-tuned CodeT5 model.
@@
Figure 4: Compilation, pass, and optimization rate per RL step for the generated programs using the fine-tuned CodeT5 model. All the rates are calculated by the number of compiled, passed, or optimized generated programs over the total number of the generated programs.
5.6 Discussion
@@
Our framework shows that an RL-based strategy for fine-tuning is able to tune a model more effectively compared to simple fine-tuning. Especially, for source code optimization, it is crucial for the model to have the ability to explore the search space by itself. However, codes that are functionally equivalent are hard to learn if their semantics are drastically different for LLM tokenizers to understand the structure of the code on limited data. Our approach encourages LLMs to make minor modifications to the input source code. However, we believe ML techniques with structure information of code can improve the ability of LLMs to understand the complicated structure difference between two source codes.
6 Conclusion
@@
In this paper, we propose a novel framework for the task of code optimization, called PerfRL. Our framework combines techniques from LLMs and RL, and it allows language models to take into consideration feedback from unit tests during their fine-tuning process. To demonstrate the applicability of our framework, we fine-tuned the CodeT5 model on the PIE dataset. We benchmark the proposed approach against a list of baseline models that rely solely on simple fine-tuning, while ignoring logical and syntactical correctness of the generated code. The evaluation results demonstrate that by adopting our framework, one can reach state-of-art performance using a smaller language model with fewer parameters, which in turn results in lower energy consumption, something that is critical on edge computing devices.
Future work directions include (i) the integration of graph representations for code such as ProGraML Cummins et al. (2020), (ii) the combination of graph neural networks and LLMs to capture both structural and language characteristics of the different parts of code, (iii) the implementation of more sophisticated scoring and reward functions for RL, (iv) the investigation of the applicability of the suggested approach to edge computing systems with strict power or energy budgets (v) the evaluation of our framework using models with different architectures but with a similar number of parameters.

### v2

@@ V Experiments
Dataset. To fine-tune and evaluate PerfRL, we use the dataset from PIE [16]. This dataset consists of approximately 40k Python files, 88k C++ files, and 3.6k Java files. PIE captures the progressive changes made by a programmer over time to improve their code. The dataset also contains (slow, fast) pairs of code written by the same programmer. We run our experiments on the subset of the dataset that concerns the Python files. The training, test and validation sets consist of approximately 36k, 1k and 2k samples respectively. Each of the samples has at least one associated unit test file that requires a specific input and has an expected output result. We test the accuracy of the dataset by executing all the input source code with a 5-second execution limit. We found that the $72.4\%$ of the training data, $76.4\%$ of the testing data, and $70.8\%$ of the validation data are executable. During our training, we skip all the input code that is not originally executable in the first place. As for testing and validation, we use all the data in order to compare our approach with the baseline model.
@@
Setup. We run the fine-tuning and reinforcement learning steps for an instance of the CodeT5 model with $60$ million parameters and a learning rate of $2\times 10^{-5}$. To reduce the training time we run the whole process for $8$ RL steps for $3$ epochs. The aforementioned RL steps and epochs have been determined by performing hyperparameter tuning. For the training of the model, we use an NVIDIA A100 GPU graphics card with 40GB of RAM on an Ubuntu 20.04 server with 50 cores. We decided to run our experiments on real hardware, as opposed to simulators such as gem5. This enables us to collect real-world evaluation results. Our model is trained within approximately 30 hours. Each epoch computes (i) the score of the generated sequence (see Equation 9) and (ii) the loss for all the samples of the dataset from the same input data. During the training phase, we set the temperature to 1 and top_k to 50 for random sampling. The source code and data to reproduce the experiments can be found in the anonymized repository of the project [^1] https://anonymous.4open.science/r/Perf_LLM-6B27/.
@@
As already mentioned, prior to running the reinforcement learning step, we fine-tune the CodeT5 model with a one-shot learning setting. That is to say, we feed each sample of the dataset into the model once and compute the loss as described in Equation 3. We set the learning rate to $5\times 10^{-5}$ and batch size to $32$. The results are available in Table LABEL:tab:results.
@@
TABLE I: Evaluation results of PerfRL and baseline models. Bold font indicates the best score on each column. The asterisk (*) indicates the baseline model. The first block shows the results as reported in [16].
@@
METHOD Model Size Sample strategy %OPT $\uparrow$ SP $\uparrow$ RTR $\uparrow$ Energy(kJ) $\downarrow$ Power(W) $\downarrow$
CODEGEN-16B 16B greedy and 1-shot 2.2 1.55 35.48 311040 1200
CODEGEN-2B 2B greedy 8.2 2.32 56.90 51840 600
CODEGEN-16B 16B greedy 14.6 1.69 40.83 311040 1200
CODEX - greedy 14.3 2.7 62.96 - -
CodeT5 (Before RL) 60M greedy and 0-shot 0 0 0 - 250
CodeT5 (24 Fine-tuning epochs)* 60M greedy and 0-shot 0.5 2.27 55.95 27000 250
PerfRL (8 RL steps, our approach) 60M greedy and 0-shot 2.8 4.93 79.72 27000 250
@@
For our experiments, we set the reward values for the reward function $r$ to $R1=0$, $R2=1$, $R3=1.3$ and $R4=2$. As already mentioned, the actual reward values do not affect the performance of the RL step, as long as we assign a higher value to a much preferable sample.
Baselines. We use a fine-tuned version of the CodeT5 model with $24$ epochs as our main baseline. Furthermore, we compare our results with those of the original PIE paper, where models from the family of CodeGen and Codex are used for their experiments.
Evaluation metrics. To evaluate our approach we use the following evaluation metrics:
  • Percent Optimized per (%OPT): The ratio of samples on the test set that are improved by a given method.
@@
  • Energy: The total energy estimation for model training.
  • Power: The total maximum power for the GPU during training.
@@
To measure the execution time of each generated program, we calculate the cumulative execution time using all the unit tests. We run each experiment three times and measure the mean time on each set of unit tests in order to ensure that the reduction of the execution time is not random.
@@
Evaluation Results. Table LABEL:tab:results presents the mean %OPT, SP, RTR, Energy and Power scores of the proposed method as well as those of the benchmark models. Although PerfRL operates with a model that has fewer parameters, it achieves performance that is equal to or even surpasses that of its counterparts. Specifically, it significantly outperforms others in terms of the RTR and SP scores with minimum energy consumption among all models. Furthermore, as illustrated in the left part of Figure 2, both the pass and optimization rate increase proportional to the RL steps. Hence, we conclude that the utilization of an RL step enables smaller models to learn easily how to optimize source code in a less resource and time-consuming manner. More specifically, we train our model for 30 hours as opposed to the baseline CodeGen which is trained for approximately 1 or 3 days and requires a stronger machine (i.e. 2 $\times$ NVIDIA A6000 for CodeGen-2B for 1 day and 4 $\times$ NVIDIA A6000 for CodeGen-16B for 3 days). Since the energy consumption is half of the total energy of CODEGEN-2B and 15 times less than CODEGEN-16B, we conclude that RL allows models to be trained more efficiently with less energy without the need for powerful machines.
@@
In order to determine whether our RL step is able to generate good candidates by random or greedy sampling during the training, we measure the compilation, pass, and optimization rate as shown in the right part of Figure 2. Considering our sampling strategy that concatenates the target program with the generated samples, we empirically observe that the threshold for these three values is approximately $20\%$ (red dashed line). Since these three rates are always above the threshold, they denote that our model generates meaningful candidates by itself other than purely based on the target program.
@@
Figure 2: (left) Pass rate and optimization rate on validation data over RL steps on the fine-tuned CodeT5 model. (right) Compilation, pass, and optimization rate per RL step for the generated programs using the fine-tuned CodeT5 model. All the rates are calculated by the number of compiled, passed, or optimized programs over the total number of the generated programs.
@@
Discussion and ablation study. Our framework shows that an RL-based strategy for fine-tuning is able to tune a model more effectively compared to simple fine-tuning. Especially, for source code optimization, it is crucial for the model to have the ability to explore the search space by itself. However, codes that are functionally equivalent are hard to learn if their semantics are drastically different for LLM tokenizers to understand the structure of the code on limited data. Our approach encourages SLMs and LLMs to make minor modifications to the input source code. However, we believe ML techniques with structure information of code can improve the ability of LLMs to understand the complicated structure difference between two source codes. Demonstrating our smaller model’s ability to compare with the larger model, such as CODEGEN-16B and CODEX and achieving better $SP$ and $RTR$ scores, serves as solid evidence of compatibility with both small and large models. Furthermore, our comparison of the CodeT5 model, using identical training steps, yielded better results than solely performing fine-tuning.
VI Conclusion
@@
In this paper, we propose a novel framework for the task of code optimization, called PerfRL. Our framework combines techniques from SLMs and RL, and it allows language models to take into consideration feedback from unit tests during their fine-tuning process. To demonstrate the applicability of our framework, we fine-tuned the CodeT5 model on the PIE dataset. We benchmark the proposed approach against a list of baseline models that rely solely on simple fine-tuning, while ignoring the logical and syntactical correctness of the generated code. The evaluation results demonstrate that by adopting our framework, one can reach state-of-art performance using an SLM with fewer parameters, which in turn results in lower energy consumption, something that is critical on edge computing devices and low-resource environments.